find the equation of a line for this points (4,-2) and (0,3)



















0

Answer by checkley71(8403)

SLOPE IS
(3+2)/(0-4)
5/-4 OR -5/4 NOW USING ONE SET OF (X,Y) POINTS IN THE LINE FORMULA WE GET THE Y INTERCEPT
Y=mX+b
3=-5/4*0+b
3=b OR
Y=-5/4X+3